# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v1

@@ Abstract
Approaching the challenge to understand fundamental physical constants (FPCs) started with understanding the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss first and second-order phase transitions and emerging new states including long-lived metastable states. We also discuss transitions between metastable states, chemical reactions and their products. These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge. This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
@@ 1 Introduction
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a) and velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b). This has deepened our understanding of these properties and their physical models.
@@ 2 Why filling the condensed matter gap is important
In the next section, we discuss several examples where FPCs set observability and operation of condensed matter effects and phenomena.
3 Observability of phase transitions and existence of different states
@@ 3.2 Second-order phase transitions
Our first case study are the second-order phase structural transitions in solids. We consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. For simplicity, we consider two phases only, quartz and $\beta$-cristobalite, its high-temperature (around 1800 K) phase Hammonds et al. (1996); Dove (1997).
@@
Compared to quartz, cristobalite has a more symmetric and open structure and a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy (5) as a result.
@@
At room temperature or any other temperature below the phase transition, cristobalite is a long-lived metastable phase of silica. Lets compare the entropy of quartz with the number of atoms $N$, $S_{Q}$, and the entropy of quartz-cristobalite mixture, $S_{Q-C}$, consisting of any proportions of the two phases and having the same total number of atoms $N$. Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined and used in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system (or system sizes small enough), the system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
$S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence, at room or other fixed temperature where both phases exist, we have
@@
We therefore find that the increase of the phase space and entropy (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase. Without a phase transition operating in the considered range, the entropy will increase according to Eq. (5) within the single quartz phase but without the frequency softening and the associated increase of the phase space (we do not consider softening of phonon frequencies with temperature due to usual anharnomicity).
We now observe that if we consider different values of FPCs, the phase transition would not operate. We recall that the upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants ($\omega_{m}$ is sometimes referred to as the “rattling” frequency) Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
@@
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (10)-(11). Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain the rest of essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
@@
Figure 2 illustrates this point for solid minerals including quartz considered in this case study. A binary quartz-cristobalite system considered with increased entropy (8) is readily generalised to a collection of many systems which exist in long-lived metastable states (see Section 3.5 for a detailed discussion of metastable states). Note that softening and the associated increase of the phase space in high-temperature phases is generally expected because these phases tend to be more symmetric and have a larger number of soft phonons as a result Hammonds et al. (1996); Dove (1997).
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. The mixture of quartz and metastable cristobalite at the same temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms). Phase transitions between different phases enable collections of systems with increased phase space and entropy at the same temperature. Shown minerals are for illustration only.
This case study gives us several insights. First, FPCs govern important effects in condensed matter and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
@@ 3.3 Superconducting transition
In the Migdal-Eliashberg theory, the upper limit to $T_{c}$ can be written as $T_{c}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (11) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum at a certain $\lambda$ because $\lambda$ larger typically above 1-2 result in instabilities and associated lowering of $T_{c}$. Using $C$ in the range 0.1-0.2 indicated by experiments, theory and modelling and Eq. (11) gives the upper bound on $T_{c}$, $T_{c}^{u}$, as Trachenko et al.
@@
Using Eq. (13), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved at our environmental conditions.
@@ 3.4 First-order phase transitions: solids, liquids, gases
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) set by FPCs in Eq. (4) because microscopically, melting involves breaking chemical bonds. Experimentally, melting temperatures correlate with cohesive energies and are on the order of $\omega_{\rm D}$ in the range 102-103 K for many systems Ubbelohde (1978); Grimvall and Sjödin (1974). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs in Eq. (10). This then gives the following role of FPCs in setting the states of matter at environmental conditions.
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only. As in the case of the second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The overlap of our environmental conditions of pressure and temperature with melting and boiling temperatures set by FPCs gives rise to two states important for life: at our environmental conditions, H2O can exist as water and carbon-based substances exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only at environmental conditions and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures substances would not exist as cohesive states but as the gas state only.
@@
$S_{g}$ in Eq. (16) is more sensitive to the variation of FPCs as compared to $S_{s}$ in Eq. (14) so that larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster (note that larger $\hbar$ and smaller $m_{e}$ and $e$ decrease the cohesive energy in Eq. (4), favouring gases over solids). Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective, however this may change if the same solid can exist as a mixture of the stable and many long-lived metastable states (see Section 3.2 and next section for a more detailed discussion).
@@ 3.5 Metastable states and chemical reactions
In three previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions and phases assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the minimum of the free energy. Nevertheless, metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition point corresponds to the equality of the free energies of both phases (see Ref. Brazhkin (2024) for review). When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2024).
As in the case of quartz-cristobalite mixture discussed in Section 3.2, the ability of a given substance to exist in many long-lived metastable states increases the phase space and entropy. Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
Transitions between metastable states involves the same processes as transitions between equilibrium states. These processes are governed by the equality of the free energy of long-lived metastable states at the transition point. Hence, similarly to what we find for equilibrium transitions in earlier sections, the observability of transitions betweeen different metastable states and hence the existence of numerous metastable phases is set by FPCs. Similarly to what we saw in Section 3.2, the phase space and entropy at a given temperature increase due to many long-lived metastable states emerging as a result of these transitions.
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and symmetry Zee (2007). FPCs would then be calculable from this theory.
In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the phase space and entropy. Metastable states were important in this discussion because at given environmental parameters (e.g., temperature), the collection of many long-lived metastable states (as opposed to just one thermodynamically stable phase) give the increase of the phase space and entropy. We discussed numerous metastable states in condensed matter physics and chemistry in Section 3. However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales of matter in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the current age of the Universe Brazhkin (2024)). These metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading new states such as the nuclear synthesis in stars).
@@
One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy. If most of the current entropy is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion of smaller BHs).
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe, identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and refining our models setting constraints on FPCs Adams (2019). Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.

### v2

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss first and second-order phase transitions and emerging new states including long-lived metastable states. We also discuss transitions between metastable states, chemical reactions and their products. These examples show that different values of FPCs would disable these transitions, and new phases as a result of these transitions would not emerge. This suggests that the current values of FPCs, by enabling transitions which give rise to new states, act to increase the phase space and entropy.
@@ 1 Introduction
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al. . This has deepened our understanding of these properties and their physical models.
@@ 2 Why filling the condensed matter gap is important
In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase and other transitions. Existence of different phases and states
@@ 3.2 Second-order phase transitions
Our first case study are the second-order phase structural transitions in solids. We consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. Let’s first consider $\beta$-cristobalite, the high-temperature phase emerging at the phase transition around 1800 K Hammonds et al. (1996); Dove (1997).
@@
As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
@@
Cristobalite is a long-lived metastable phase of silica which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times $\Delta t$ shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties agree with those described by equilibrium thermodynamic functions Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
As a more general case compared to pure metastable cristobalite, let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below the transition point where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
@@
We therefore find that the entropy increase (8) is ultimately related to a phase transition which gives rise to the second higher-entropy phase. The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state at low $T$.
Without a phase transition operating in the considered range, the entropy would increase according to Eq. (5) within the single quartz phase but without the higher entropy contribution from the second metastable phase.
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ in condensed matter can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ as
@@
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (10)-(11). Note that this question is well-posed because it is possible to change $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
@@
A binary quartz-cristobalite mixture with increased entropy (8) is readily generalised to a collection of many systems existing in long-lived metastable states. Figure 2 illustrates this point. The increase of the phase space and entropy in these metastable states is generally expected because (a) high-temperature phases tend to be more symmetric Landau and Lifshitz (1970) with more open structure and have a larger number of soft phonons as a result Hammonds et al. (1996); Dove (1997) and (b) the increased phase space and entropy in the high-temperature phases become “frozen-in” in metastable states at low $T$.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. The mixture of quartz and metastable cristobalite at low temperature has combined phase space and entropy that are larger than those of quartz alone (with the same number of atoms). Phase transitions between different states enable sets of long-lived metastable systems with higher combined entropy at the same temperature. Shown minerals are for illustration only.
This case study gives us several insights. First, FPCs govern key effects in condensed matter physics and define the regime in which the system is, classical or quantum. Second, FPCs affect the observability and operation of entire effects such as phase transitions. Third, FPCs, by enabling phase transitions, can increase the phase space and entropy.
@@ 3.3 Superconducting transition
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (11) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (11), this gives $T_{c}^{u}$ as Trachenko et al.
@@
Using Eq. (13), we ask how would the observability of superconductivity change were fundamental constants to take different values? Since $T_{c}^{u}$ is the upper bound, $T_{c}^{u}$ on the order of 10-5 K or lower due to different FPCs implies that superconductivity would be unobserved.
@@ 3.4 First-order phase transitions: solids, liquids, gases
The triple point pressure of many substances is below the atmospheric pressure, and so we often observe melting and boiling transitions on temperature increase. The phase co-existence lines in Figure 3 are defined by the equality of thermodynamic potentials. At high pressure, melting temperature substantially depends on pressure (see Ref. Trachenko (2024) for the functional form of pressure-temperature melting lines). At low pressure close to the triple point, melting can be discussed on the basis of cohesive energy Ubbelohde (1978) because microscopically, melting involves breaking chemical bonds. This cohesive energy is set by FPCs in Eq. (4). Melting temperatures correlate with cohesive energies and are on the order of $\omega_{m}$ or $\omega_{\rm D}$ in the range 102-103 K in many systems Ubbelohde (1978); Grimvall and Sjödin (1974) (the difference between the cohesive energy and $\hbar\omega_{m}$ is due to the small factor $\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$ in Eq. (9)). As discussed in Section 3.2, $\omega_{m}$ and $\omega_{\rm D}$ are set by FPCs. This gives the following role of FPCs in setting the states of matter.
FPCs resulting in large cohesive energies and melting temperatures imply that matter would exist in the solid state only at given environmental conditions. Melting and boiling phase transitions would be unobservable. Conversely, FPCs resulting in low cohesive energies and melting temperatures imply that matter would exist in the gas state only.
We see that, similarly to second-order phase transitions discussed in Section 3.2, the current values of FPCs set the observability of first-order phase transitions in Figure 3 and the existence of different states of matter.
The apparent overlap of our environmental conditions and melting and boiling temperatures set by FPCs gives rise to two states important for life. First, H2O can exist as liquid water at our typical planetary conditions. Second, carbon-based substances can exist as bonded solidlike structures. Much larger cohesive energy due to different values of FPCs would imply that H2O could exist as a solid only and hence no water-based life. Much smaller values of cohesive energies due to different FPCs would imply that H2O and carbon-based structures would not exist as cohesive states but as the gas states only.
@@
The same FPCs interestingly feature in both numerators and denominators in Eqs. (14) and (16). $S_{g}$ in Eq. (16) is more sensitive to FPCs as compared to $S_{s}$ in Eq. (14): larger $\hbar$ and smaller $m_{e}$ and $e$ increase $S_{g}$ faster. The same variation - larger $\hbar$ and smaller $m_{e}$ and $e$ also decrease the cohesive energy in Eq. (4), favouring gases over solids. Hence gases may be viewed as more entropically favourable over solids from the FPCs perspective.
@@ 3.5 Metastable states and chemical reactions
In previous sections, we discussed the observability of phase transitions and the existence of different states as a result of these phase transitions. The usual statistical-mechanical treatment of these transitions assumes equilibrium phases corresponding to the global minimum of the free energy. However, many simple molecular compounds of light elements (e.g., most hydrocarbons, nitrogen oxides, hydrates, carbides, carbon oxide, alcohols, glycerin) are metastable and do not correspond to the lowest minimum of the free energy. Nevertheless, long-lived metastable phases in a certain range of pressure and temperature undergo reversible phase transitions which are consistent with laws of equilibrium thermodynamics: the phase transition corresponds to the free energy crossover (see Ref. Brazhkin (2006) for review). When discussing phase transitions and phase diagrams in these compounds, the tacit assumption is the one we discussed in Section 3.2: the observation time is assumed to be shorter than the relaxation time of the long-lived metastable state.
Going from elemental and simple substances to more complex compounds including organic compounds, we find that nearly all of them are metastable and do not correspond to the lowest free energy minimum for a given composition. The world of organic compounds relevant for life processes is nearly all metastable Brazhkin (2006). Considering large inorganic structures suggests that the associated phase space can be astronomical if all possible conformations are considered (not accounting for the smaller number of actual viable existing structures) Cortes et al. .
Transitions between metastable states involve the same atomistic processes as transitions between equilibrium states and energies related to the chemical bond energy set by FPCs. As in the equilibrium case, these transitions correspond to the free energy crossover. Hence, similarly to what we saw for equilibrium transitions in earlier sections, the observability and operation of transitions between different metastable states and hence the existence of numerous metastable phases is set by FPCs.
Similarly to what we saw in Section 3.2, the entropy increases due to new metastable states emerging as a result of these transitions because the states formed at higher temperature tend to be more symmetric with more open structure and have larger phase space as a result (this larger phase space and higher entropy become frozen in the low-temperature metastable phase).
@@ 4 Fundamental constants and entropy increase
We consider that understanding the current values of FPCs requires the discussion of their variation Barrow (2003); Adams (2019). A different perspective should be noted: a final theory of most fundamental laws of nature, when found, would be rigid (“logically isolated”) Weinberg (1993) and guided by a unique design and an exquisite symmetry Zee (2007). FPCs would then be calculable from this theory.
In previous sections, we discussed examples in condensed matter physics where FPCs, by enabling phase transitions and different phases, increase the entropy. Long-lived metastable states were important in this discussion because at given environmental parameters, the collection of these states (as opposed to just one single thermodynamically stable phase with the global minimum) increase the phase space and entropy. However, as recently reviewed Brazhkin (2024), metastable states are ubiquitous at different levels and scales in the Universe including nuclei and astronomical objects (energy barriers setting the life time of these states can vary in a wide range 10-7 to 109 eV per particle, with corresponding relaxation times often exceeding the age of the Universe Brazhkin (2024)). These long-lived metastable states, similarly to those discussed in Section 3, (a) emerge as a result of different transitions and reactions which are sensitive to the values of fundamental constants Barrow (2003); Adams (2019) and (b) can co-exist at the same environmental parameters (these parameters can be themselves affected by reactions leading to new states such as the nuclear synthesis in stars. In this case, entropy change in the environment due to heat release needs to be taken into account in the entropy balance).
@@
One example illustrating some of the discussion in the previous paragraph, but which needs further study, is the black hole (BH) entropy. If most of the current entropy in the Universe is associated with the entropy of large BHs Egan and Lineweaver (2010), one might consider small $\hbar$ and $c$ to maximise the BH entropy
@@
where $M_{BH}$ is the BH mass. However, small $\hbar$ and $c$ give large $\alpha$ in Eq. (2), disrupting many key processes of structure formation at stages preceding the formation of BHs, including stellar evolution which affect BHs at later times Barrow (2003); Adams (2019) (the mechanism of formation of large BHs is currently under discussion and includes the accretion involving smaller BHs).
This discussion and the conjecture of FPCs acting to increase the entropy need further study. This includes detailing entropy changes in processes at different levels of structure organisation in the Universe Barrow (2003); Adams (2019), identifying sets of subsystems which can be considered closed and where the entropy argument can be applied, clarifying the role of gravity in this process Landau and Lifshitz (1970) and so on. Given the enormity of the problem we face in understanding the origin of fundamental constants Barrow (2003); Weinberg (1983); Barrow and Webb (2006), making new observations and asking new questions is useful for progress.